### v1

# Binary Cumulative Encoding meets Time Series Forecasting

Abstract
Recent studies in time series forecasting have explored formulating regression via classification task. By discretizing the continuous target space into bins and predicting over a fixed set of classes, these approaches benefit from stable training, robust uncertainty modeling, and compatibility with modern deep learning architectures. However, most existing methods rely on one-hot encoding that ignores the inherent ordinal structure of the underlying values. As a result, they fail to provide information about the relative distance between predicted and true values during training. In this paper, we propose to address this limitation by introducing binary cumulative encoding (BCE), that represents scalar targets into monotonic binary vectors. This encoding implicitly preserves order and magnitude information, allowing the model to learn distance-aware representations while still operating within a classification framework. We propose a convolutional neural network architecture specifically designed for BCE, incorporating residual and dilated convolutions to enable fast and expressive temporal modeling. Through extensive experiments on benchmark forecasting datasets, we show that our approach outperforms widely used methods in both point and probabilistic forecasting, while requiring fewer parameters and enabling faster training.
1 Introduction
Time series forecasting plays a key role in decision-making across domains such as finance [akita2016deep], healthcare [song2020distributed], retail [bandara2019sales], and climate science [zaini2022systematic]. While classical methods like ARIMA and Exponential Smoothing remain widely used in practice [hyndman2018forecasting], recent advances increasingly favor deep learning approaches [lim2021time, li2024deep], particularly in scenarios requiring high-precision forecasting [zhang2024probts].
A critical component of deep learning-based forecasting is data preprocessing. This typically includes normalization, min-max or mean scaling [salinas2020deepar], Box-Cox transformation [hyndman2018forecasting], and short-time Fourier transforms [furman2023short]. Recent studies [ansari2024chronos, liu2024timer] have explored quantization, discretizing the continuous target space into fixed-size bins. Values are then represented as one-hot vectors, enabling a tokenized view of time series aligned with Transformer-style architectures. Empirical results suggest that such representations enhance performance, especially within Time Series Foundation Models [zhang2024probts].
However, one-hot encoding lacks ordinal awareness, treating bins as unrelated categories and ignoring relative distances–important cues for modeling temporal dynamics. To address this, we propose Binary Cumulative Encoding (BCE): a representation where all bins below (or equal to) the current value are activated. This monotonic encoding, previously applied to tabular data [gorishniy2022embeddings], preserves global ordinal structure and yields richer inputs for neural models. Figure 1 shows an example of BCE applied to a time series.
Figure 1: Visualization of different representations of time series data. (left) Original time series showing monthly airline passenger counts. (center) One-hot encoding of discretized values. (right) Binary Cumulative Encoding (BCE), where all bins less than or equal to the active bin are highlighted, preserving ordinal structure.
Choosing an appropriate neural architecture is another central design decision. RNNs, CNNs, and Transformers are commonly used, with Transformer-based models [zhang2024probts] excelling due to their capacity for long-range temporal modeling [vaswani2017attention]. Yet, their nearly quadratic complexity with sequence length [tay2022efficient] poses practical challenges.
To overcome this, we introduce a CNN-based model that combines 2D and 1D convolutions to process BCE representations efficiently. This design reduces computational cost while preserving representational expressiveness. Moreover, BCE naturally supports probabilistic output by modeling distributions over bins–an ability often missing in standard CNN forecasting models.
We validate our method on the ProbTS benchmark [zhang2024probts], which includes diverse datasets from multiple domains. Our model consistently outperforms strong baselines in both point and probabilistic forecasting tasks.
The main contributions of this work are:
  • A novel discrete representation, Binary Cumulative Encoding (BCE), which captures ordinal and metric structure in time series.
  • A lightweight convolutional model, BinConv, combining 2D and 1D convolutions to process BCE sequences efficiently.
  • Extensive evaluation on the ProbTS benchmark, showing that our method achieves state-of-the-art results across multiple forecasting datasets.
2 Related Work
Classical and Deep Learning Methods for Time Series Forecasting. Classical models such as ARIMA, exponential smoothing, and vector autoregression are widely used for their simplicity and efficiency [hyndman2018forecasting]. However, they often underperform on large-scale datasets exhibiting complex and nonlinear patterns, as shown in M4, M5, M6 forecasting competitions [makridakis2018m4, makridakis2022m5, makridakis2024m6]. Deep learning approaches improve upon these by framing forecasting as a sequence-to-sequence problem using fixed context windows [lim2021time]. Recent developments like Time Series Foundation Models [liang2024foundation, das2024decoder, darlow2024dam] leverage large-scale pretraining to generalize across domains, achieving strong results in both point and probabilistic forecasting [zhang2024probts].
Neural Network Architectures for Time Series Forecasting. Neural models vary in their trade-offs between accuracy, scalability, and speed. Transformer-based models such as PatchTST [nie2023a], TFT [lim2021temporal], and Informer [zhou2021informer] capture long-range dependencies well, but incur high computational costs due to their quadratic scaling with sequence length. Earlier RNN-based models like LSTM [hochreiter1997long], GRU [che2018recurrent], and DeepAR [salinas2020deepar] are more efficient at inference, though limited by their sequential nature, which hinders training parallelism. CNN-based architectures such as TCN [bai2018empirical], DeepTCN [chen2020probabilistic], and ModernTCN [luo2024moderntcn] offer fast, parallel training and strong forecasting performance. Following this line, our CNN-based BinConv model balances accuracy and efficiency when used with BCE inputs.
Token-Based Representations of Time Series. Inspired by NLP, recent work reformulates time series forecasting as next-token prediction using discretization. Models like LLMTime [gruver2023large], TiMER [liu2024timer], and Chronos [ansari2024chronos] encode values as digits, patches, or bins. Gorishniy et al. [gorishniy2022embeddings] proposed piecewise linear encoding (PLE) for tabular data, preserving ordinal structure. Most models rely on one-hot encoding, which discards distance and order information between bins. This limits the ability to learn value relationships. In contrast, our BCE encoding captures both order and magnitude, improving learning while remaining compatible with token-based models.
Probabilistic Forecasting with Deep Learning Methods. Probabilistic forecasting offers both predictions and uncertainty. Deep learning models approach this via quantile regression (e.g., TFT, DeepAR), parametric output distributions (e.g., DeepTCN), or generative models like TimeGrad [rasul2021autoregressive] and GRU-NVP [rasul2021multivariate]. While effective, generative approaches can be unstable and costly to train. Token-based models such as Chronos and TiMER enable probabilistic inference via sampling in discrete space. Our model adopts this approach using BCE, supporting both point and probabilistic forecasts through structured sampling.
3 Binary Cumulative Encoding
[…]
Figure 2: Example of BCE encoding for 4 bins.
The scale of time series can vary significantly, even within a single dataset. Therefore, before converting time series values into an encoding space, we apply normalization to map values into a quantized range. Similar to the Chronos model [ansari2024chronos], we use the mean scaling technique [salinas2020deepar], where each value is divided by the mean over a given context window. Specifically, we compute the scaled value as $\tilde{x}=x/s$, where $s=\frac{1}{C}\sum_{i=1}^{C}\left|x_{i}\right|$.
Next, we quantize the scaled time series as follows:
$\mathrm{BCE}(x)$ $=[v_{1},\dots,v_{D}]\in\mathbb{R}^{D}$
$v_{d}$ =$=\begin{cases}0,&\tilde{x}<b_{d}\\ 1,&\tilde{x}\geq b_{d}\end{cases},$ (1)
where $b_{d}$ indicates a $d$-th bin in the original space. An example of construction of BCE is shown in Figure 2.
The interval from $b_{0}$ to $b_{D}$ is uniformly split into $D$ bins in a data-agnostic manner, where $D$, $b_{0}$, and $b_{D}$ are hyperparameters. In our experiments, we do not fine-tune these values per dataset and use fixed settings of $D=1000$, $b_{0}=-5$, and $b_{D}=5$, unless stated otherwise. This quantization results in binary cumulative encoding of the form $111\ldots 100\ldots 0$.
This type of encoding is not new in the context of tabular deep learning for numerical features [gorishniy2022embeddings]. However, unlike previous work, we not only transform the inputs using this scheme, but also forecast targets directly in the encoding space, as detailed in Section 4.2. After generating a forecast in the encoding space, we apply an inverse transformation by assigning the average value of the bin encoding, i.e., $\tilde{y}=(b_{i}+b_{i+1})/2$, where $i$ is the index of the last $1$. Finally, to obtain the forecast in the original space, we reverse the mean scaling: $y=s\cdot\tilde{y}$. To the best of our knowledge, this approach is novel in the context of time series forecasting.
4 BinConv
4.1 Architecture
Input $\mathbb{R}^{C}$ BCE $\mathbb{R}^{C\times D}$ $M$ BinConv block $\mathbb{R}^{C\times D}$ Conv1d Sigmoid $\mathbb{R}^{D}$ Inverse BCE $\mathbb{R}$
(a) BinConv architecture
Input $\mathbb{R}^{C\times D}$ Conv2d DyTanh $\mathbb{R}^{K\times D}$ Conv1d-1 ReLU $\mathbb{R}^{K\times D}$ Conv1d-2 ReLU Dropout $\mathbb{R}^{C\times D}$ + Output $\mathbb{R}^{C\times D}$
(b) BinConv block
Figure 3: BinConv architecture. (a) First, we preprocess the time series using BCE. Then, the BinConv model is trained in the BCE space, and during forecasting, the inverse BCE transformation is applied to obtain predictions in the original space. BinConv consists of $M$ stacked BinConv blocks followed by a final 1D convolution layer with a large kernel size. (b) Each BinConv block is the core component of the architecture, comprising three convolutional layers and a residual connection. The DyTanh activation function is used as a substitute for layer normalization. The output dimension of each BinConv block matches its input dimension.
Table 1: Convolutional parameters used in BinConv.
Layer Kernel Size In/Out Channels Groups
Conv2d $(C,\;s_{1})$ $1\ /\ K$ $1$
Conv1d-1 $s_{2}$ $K\ /\ K$ $K$
Conv1d-2 $s_{2}$ $K\ /\ C$ $K$
[…]
To model time series effectively in the discrete BCE space, the architecture must preserve ordinal relationships, support autoregressive forecasting, and maintain computational efficiency. To this end, we propose BinConv, a convolutional neural architecture that efficiently captures the local temporal and monotonic structure of BCE vectors by combining 2D and 1D convolutions.
The BinConv model consists of $M$ convolutional blocks followed by a final convolutional layer with a sigmoid activation function, producing an output vector in the BCE space (see Figure 3(a)). For each convolution, we add padding on both sides to preserve output dimensionality. To preserve the cumulative structure of the BCE representation, we use ones for left padding and zeros for right padding. Notably, we do not use fully connected layers in the model. In fact, including fully connected layers degrades performance, as further discussed in Section 6.1.
Each BinConv block comprises three convolutional layers (see Figure 3(b)). The first is a 2D convolution with kernel size $(C,3)$, where $C$ corresponds to the context length. This layer transforms the input matrix–obtained after applying BCE–into $K$ one-dimensional vector representations, where $K$ is the number of output channels and treated as a hyperparameter. Following the work of [zhu2025transformers], we use the DyTanh activation function, which replaces layer normalization in our architecture.
Next, we apply two 1D convolutional layers with ReLU activation, followed by dropout after the second layer. Inspired by TCN architectures [bai2018empirical, luo2024moderntcn], we employ depthwise convolution with the number of groups set to $K$, which significantly reduces computational cost without sacrificing performance 6.2. Each BinConv block is designed such that the input and output dimensions are identical. This allows the use of a residual connection as the final operation in the block, enabling the stacking of multiple BinConv blocks while mitigating the vanishing gradient problem. The main parameters of the model are listed in Table 1.
To train the model, we apply the binary cross-entropy loss to the model output. This approach is analogous to a multi-label classification setup, where each class is predicted independently using a "one-versus-all" scheme.
4.2 Forecasting
To forecast time series, we apply an autoregressive approach: forecasting at time step $t+1$, appending the result to the context, then forecasting at $t+2$, and so on. To do this, we need to transform the vector of sigmoid-activated outputs into a BCE. We interpret each component of the output as the success probability of an independent Bernoulli trial. Due to the design of the BCE, only sequences of the form $11\ldots 10\ldots 0$ are considered valid. Therefore, we normalize probability of each valid sequence by the total probability of all valid sequences, denoted as $Z$.
More formally, the probability of a valid encoding vector $e_{m}$, where the first $m$ components are 1 and the remaining $D-m$ components are 0, is given by:
$p(e_{m})=\frac{\hat{p}(e_{m})}{Z}=\frac{\left(\prod_{i=1}^{m}p_{i}\right)\left(\prod_{i=m+1}^{D}(1-p_{i})\right)}{\sum_{j=0}^{D}\left(\prod_{i=1}^{j}p_{i}\prod_{i=j+1}^{D}(1-p_{i})\right)},$ (2)
where $p_{i}$ is the sigmoid output at position $i$.
Given these probabilities, we can either sample $n$ trajectories autoregressively to obtain a probabilistic forecast, or take the argmax at each time step, which is faster but results in a point forecast. In Section 5, we show that the point estimates of the point and probabilistic forecasts do not differ significantly, and both outperform the point estimates of other models.
[…]
$\Downarrow$
000
$\hat{p}(e_{0})=0.048$
$p(e_{0})=\frac{0.048}{Z}\approx 0.109$
100
$\hat{p}(e_{1})=0.032$
$p(e_{1})=\frac{0.032}{Z}\approx 0.073$
110
$\hat{p}(e_{2})\approx 0.288$
$p(e_{2})=\frac{0.288}{Z}\approx\mathbf{0.654}$
111
$\hat{p}(e_{3})=0.072$
$p(e_{3})=\frac{0.072}{Z}\approx 0.164$
$Z=\sum_{m}\hat{p}(e_{m})=0.048+0.032+0.288+0.072=0.44$
Figure 4: Converting the neural network output $[0.4,0.9,0.2]$ into a BCE, the most probable sequence is $[1,1,0]$.
5 Experiments
5.1 Setup
In this paper, our primary focus is on univariate forecasting. We evaluate performance on three univariate datasets: M4 Daily, M4 Weekly, and Tourism Monthly [makridakis2020m4, athanasopoulos2011tourism]. The prediction horizon $H$ for each dataset is set to $14$, $13$, and $24$, respectively. The context length is set to $C=3H$ for all datasets.
We adopt the training and evaluation setup from [zhang2024probts]. Specifically, we train each model for $50$ epochs, then select the best checkpoint based on performance on the training set, and report evaluation metrics on the test set. For all models, we used the Adam optimizer [kingma2015adam] for training. We compute CRPS using $20$ quantile levels and $100$ samples to assess the quality of probabilistic forecasts, and NMAE to evaluate point forecasting performance following to [zhang2024probts]. All computations were done on NVIDIA Tesla A10 22.5GB GPU RAM.
5.2 Results
We evaluated the performance of five different models: BinConv (our model), GRU-NVP, TimeGrad, and PatchTST, which are briefly described in Section 2. We also added simple linear model DLinear [zeng2023transformers] for the comparison. For all baseline models, we use the optimal set of hyperparameters and normalization strategies provided by [zhang2024probts]. For BinConv, as discussed in Section 3, we used mean scaling before quantization. Mean scaling was applied independently to each sample, as the M4 datasets consist of diverse time series with varying scales. We did not fine-tune the BinConv model’s hyperparameters for each dataset, as it already demonstrated strong performance with the default configuration. The hyperparameters used across all datasets are summarized in Table 2.
Table 2: Default hyperparameters used for BinConv across all datasets. $C$ denotes the context length.
Hyperparameter Value
Number of bins ($D$) $1000$
Minimum bin value ($b_{0}$) $-5$
Maximum bin value ($b_{D}$) $5$
Number of output channels ($K$) $C$
Kernel size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel size of final 1D convolution ($s_{3}$) $51$
Number of BinConv blocks ($M$) $3$
Dropout $0.35$
Learning rate $0.001$
The comparison between models is presented in Table 3. We trained and evaluated all models using five different random seeds and report the minimum, average, and maximum scores for each model. The metrics of BinConv are consistently better than those of the baseline models. Notably, the average scores of BinConv for both CRPS and NMAE are better than the best scores of all baselines.
Additionally, Table 4 compares the performance of BinConv when using argmax-based forecasting versus sampling-based forecasting (see Section 4.2 for details). While using point forecasting slightly degrades performance, it still outperforms all baseline models. We provide additional results on multivarite time-series datasets in Appendix C.
Table 3: CRPS and NMAE results for all models. Each metric is reported as the average (AVG), minimum (MIN), and maximum (MAX) over 5 different seeds.
Dataset Stat DLinear GRU-NVP PatchTST TimeGrad BinConv
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
m4_daily AVG 0.0440 0.0440 0.0714 0.0782 0.0438 0.0438 0.0587 0.0696 0.0327 0.0382
m4_weekly AVG 0.1046 0.1046 0.0972 0.1062 0.1039 0.1039 0.1046 0.1283 0.0902 0.0972
tourism_monthly AVG 0.2745 0.2745 0.3723 0.4536 0.2633 0.2633 0.2295 0.2968 0.1824 0.1955
m4_daily MIN 0.0394 0.0394 0.0570 0.0649 0.0404 0.0404 0.0402 0.0475 0.0291 0.0341
m4_weekly MIN 0.1007 0.1007 0.0916 0.1034 0.1020 0.1020 0.0800 0.0985 0.0851 0.0921
tourism_monthly MIN 0.2595 0.2595 0.2880 0.3849 0.2281 0.2281 0.1993 0.2596 0.1777 0.1911
m4_daily MAX 0.0506 0.0506 0.1196 0.1259 0.0476 0.0476 0.0933 0.1077 0.0366 0.0427
m4_weekly MAX 0.1072 0.1072 0.1006 0.1085 0.1078 0.1078 0.1718 0.2121 0.0981 0.1055
tourism_monthly MAX 0.3077 0.3077 0.5799 0.5399 0.2888 0.2888 0.2500 0.3186 0.1892 0.2012
Table 4: Comparison of BinConv performance using sampling versus argmax forecasting. Each metric is reported as the average (AVG), minimum (MIN), and maximum (MAX) over 5 different seeds.
Dataset Stat Argmax Sampling
CRPS NMAE CRPS NMAE
m4_daily AVG 0.0371 0.0371 0.0327 0.0382
m4_weekly AVG 0.0975 0.0975 0.0902 0.0972
tourism_monthly AVG 0.1961 0.1961 0.1824 0.1955
m4_daily MIN 0.0348 0.0348 0.0291 0.0341
m4_weekly MIN 0.0927 0.0927 0.0851 0.0921
tourism_monthly MIN 0.1916 0.1916 0.1777 0.1911
m4_daily MAX 0.0387 0.0387 0.0366 0.0427
m4_weekly MAX 0.1049 0.1049 0.0981 0.1055
tourism_monthly MAX 0.2013 0.2013 0.1892 0.2012
5.3 Efficiency of the Model
In Table 5, we report the total number of parameters for each model across all datasets. The DLinear model has significantly fewer parameters than the others, with BinConv being the second most parameter-efficient.
In Table 6, we present the average computation time during both training and inference for all models on the M4 Daily dataset. To obtain these results, we trained and evaluated all models using a batch size of $1$, and computed the mean and standard deviation over all samples and epochs. As expected, DLinear is the fastest model during training, with BinConv being the second fastest.
During inference, PatchTST is the fastest model because it is non-autoregressive and predicts the entire forecast horizon in a single forward pass. DLinear ranks second, followed by BinConv and GRU-NVP, which exhibit similar inference speeds–approximately twice as slow as PatchTST. TimeGrad is significantly slower than all other models.
Using point (argmax-based) forecasting instead of sampling improves inference speed by approximately 25%. However, the difference is not dramatic, as sampling is implemented by duplicating the input $N$ times, where $N$ is the number of samples. This allows all $N$ samples to be processed in a single batch.
Table 5: Total number of parameters (in thousands) per method and dataset.
Dataset DLinear BinConv PatchTST TimeGrad GRU-NVP
m4_daily $1.204$ $20.173$ $47.470$ $107.434$ $141.108$
m4_weekly $1.040$ $17.680$ $43.725$ $107.434$ $141.108$
tourism_monthly $3.504$ $54.013$ $155.288$ $107.434$ $141.108$
Table 6: Mean and standard deviation of time (in seconds) to process a single sample for the M4 Daily dataset.
Phase DLinear BinConv PatchTST TimeGrad GRU-NVP
Argmax Sampling
Train $\mathbf{0.0123\pm 0.0026}$ $0.0284\pm 0.0043$ $0.0284\pm 0.0043$ $0.0381\pm 0.0049$ $0.0450\pm 0.0063$ $0.0362\pm 0.0048$
Test $0.0755\pm 0.0016$ $0.09684\pm 0.00306$ $0.12788\pm 0.0017$ $\mathbf{0.0547\pm 0.0021}$ $2.7723\pm 0.0365$ $0.1239\pm 0.0086$
6 Ablation Study
In this section, we conduct ablation studies to better understand the contribution of core components in our model architecture. Specifically, we investigate the effect of replacing the final layer with a fully connected layer (Section 6.1), substituting Depthwise convolution layers in the BinConv block with standard convolution layers (Section 6.2), and using DyTanh instead of layer normalization (Section 6.3). Each component is modified independently, while the rest of the architecture and hyperparameters remain unchanged from the standard BinConv model. During the prediction stage, we used probabilistic forecasting to measure the difference between both probabilistic and point estimates. To save computation time, we ran all modified models using a single seed and compared their performance with the average performance of BinConv.
Table 7: Comparison between the standard BinConv model and its architectural variations. The notation w/ indicates that the standard BinConv is modified by replacing a specific component: w/ FC Layer uses a fully connected layer instead of the final convolution, w/ Standard Conv replaces Depthwise convolutions with standard ones, and w/ LayerNorm substitutes DyTanh with layer normalization.
Dataset BinConv w/ FC Layer w/ Standard Conv w/ LayerNorm
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
m4_daily $0.0327$ $0.0382$ $0.0411$ $0.0468$ $0.0292$ $0.0343$ $0.0311$ $0.0365$
m4_weekly $0.0902$ $0.0972$ $0.0870$ $0.0943$ $0.1028$ $0.1112$ $0.0847$ $0.0921$
tourism_monthly $0.1824$ $0.1955$ $0.1756$ $0.1910$ $0.1801$ $0.1945$ $0.1761$ $0.1889$
Table 8: Number of parameters (in thousands) for the standard BinConv model and its architectural variations. See the caption of Table 7 for the explanation of the column notation.
Dataset BinConv w/ FC Layer w/ Standard Conv w/ LayerNorm
m4_daily $20.173$ $1019.540$ $51.679$ $20.170$
m4_weekly $17.680$ $1017.164$ $44.830$ $17.677$
tourism_monthly $54.013$ $1052.210$ $146.899$ $54.010$
6.1 Using Linear Layer in the Output
When designing the BinConv architecture, we deliberately avoided using fully connected layers. There are two main reasons for this. First, fully connected layers introduce significantly more trainable parameters compared to convolutional layers, especially considering that the our model’s output lies in a high-dimensional space. Second, because of the design of BCE and the way we compute the loss during training, using a fully connected layer, especially when it serves as the final layer, may limit BinConv’s ability to forecast values on the test set that fall outside the range observed during training. We elaborate on this in detail in Appendix B.
Nevertheless, since fully connected layers are a popular design choice, we evaluate the performance of BinConv when a fully connected layer is used as the final layer. To implement this variant, we replace the final 1D convolutional layer as follows. The output of the BinConv blocks is a vector $h\in\mathbb{R}^{C\times D}$, where $C$ is the context length and $D$ is the number of bins. We first average $h$ over the context dimension, resulting in a vector $\bar{h}\in\mathbb{R}^{D}$. A fully connected layer is then applied to map $\mathbb{R}^{D}$ to $\mathbb{R}^{D}$.
In Table 7, we report the evaluation metrics for this variant in the column labeled “w/ FC Layer”. As we expected, the performance degrades significantly. Furthermore, as shown in Table 8, adding a fully connected layer dramatically increases the number of trainable parameters, making the model less efficient.
6.2 Using Depthwise instead of Standard Convolution
Depthwise convolution is a form of convolution in which each input channel is convolved independently with its own filter. As a result, it requires significantly fewer convolutional kernels, which improves the model’s computational efficiency. In our model, we use this type of convolution in the BinConv block for 1D operations.
The performance of BinConv with standard 1D convolutional layers in place of depthwise layers is reported in Table 7, in the column labeled “w/ Standard Conv”. Replacing depthwise convolutions with standard ones does not improve performance, but it does reduce model efficiency, as shown in Table 8.
6.3 Using Layer Normalization instead of DyTanh
In [zhu2025transformers], the authors proposed using DyTanh instead of layer normalization for transformer models. They demonstrated empirically that it yields comparable or better results in computer vision (CV) and natural language processing (NLP) tasks.
Inspired by this finding, we replace layer normalization with DyTanh in BinConv. In this section, we assess whether substituting DyTanh with traditional layer normalization followed by a ReLU activation affects performance. As shown in Table 7, the performance of this model variant is similar to or better than that of the standard BinConv, indicating that layer normalization is also a valid choice for our architecture. The number of parameters is also comparable (Table 8); DyTanh introduces only one additional parameter compared to the normalization layer, which is a negligible difference.
7 Conclusion
In this paper, we proposed using Binary Cumulative Encoding (BCE) as a discrete quantization technique for time series, and we introduced BinConv, a convolutional architecture specifically designed for efficient processing of BCE vectors. Extensive evaluation on the publicly available ProbTS benchmark demonstrates that our method achieves state-of-the-art performance across diverse domains.
Despite these promising results, our method has several limitations. First, BCE discretizes time series values, which requires approximating the precision of continuous data and can result in information loss. Second, selecting the number of bins introduces a trade-off: too few bins reduce representational fidelity, while too many increase computational complexity. Third, the current design is limited to univariate time series, which necessitates modeling each variable independently and prevents capturing cross-series interaction.
Future directions for this work include evaluating the generalization capability of our method in a Time Series Foundation Model setting, where the model is tested on previously unseen datasets. Another promising avenue is to extend the BinConv architecture to support multivariate time series by modeling interactions between components. Additionally, our approach could be applied to other time series tasks, such as imputation, anomaly detection, and classification.
[…]
Appendix A Evaluation on Additional Datasets
In this section, we report results for long-term forecasting on several datasets commonly used in the research community: ETTh1, ETTh2, ETTm1, ETTm2, and Illness [zhang2024probts]. For baseline models, we used the optimal set of hyperparameters from [zhang2024probts], while for BinConv we used the default configuration without any fine-tuning (see Table 10). The context length and prediction length were set to $36$ for the Illness dataset and $96$ for the others.
For each dataset and model, we ran only a single seed due to the computational cost of evaluation—particularly for models like CSDI and TimeGrad, which can require up to 20 hours for training and evaluation. All datasets are multivariate and contain 7 variables each. Since BinConv currently supports only univariate forecasting, and extending it to multivariate settings is left for future work, we forecasted each time series independently using the sampling procedure described in Section 4.2. Table 11 reports the total number of parameters for each model. BinConv is less parameter-efficient in the multivariate setting than in the univariate case, as it maintains a separate model for each time series in the dataset.
In Table 9, we report CRPS and NMAE for all baseline models and BinConv. As there is no clear winner across all datasets, we additionally report model rankings for both metrics in Table 12. Rankings were calculated based on relative performance: the best-performing model on each dataset received rank 1, the second-best received rank 2, and so on.
Despite the fact that BinConv’s hyperparameters were not tuned and the model is not specifically designed for multivariate forecasting, it achieved the best average rank based on the CRPS metric and the second-best rank based on the NMAE metric.
Table 9: Forecasting performance (CRPS and NMAE) across different models and datasets. All datasets use context length 96 and prediction length 96 (except Illness, it uses 36 for both).
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTh1 0.3339 0.3339 0.3218 0.3218 0.3435 0.4383 0.2930 0.3867 0.3830 0.4871 0.2969 0.3201
ETTh2 0.2319 0.2319 0.1763 0.1763 0.2623 0.3100 0.2059 0.2574 0.2291 0.3004 0.1709 0.1853
ETTm1 0.2663 0.2663 0.2681 0.2681 0.4029 0.5018 0.4201 0.5153 0.2397 0.3124 0.2922 0.3141
ETTm2 0.1431 0.1431 0.1337 0.1337 0.2106 0.2618 0.1595 0.1930 0.1116 0.1413 0.1329 0.1456
Illness 0.1869 0.1869 0.1103 0.1103 0.0750 0.0828 0.0643 0.0868 0.3077 0.3237 0.1187 0.1447
Table 10: Default hyperparameters used for BinConv across long term datasets. $C$ denotes the context length.
Hyperparameter Value
Number of bins ($D$) $500$
Minimum bin value ($b_{0}$) $-5$
Maximum bin value ($b_{D}$) $5$
Number of output channels ($K$) $C$
Kernel size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel size of final 1D convolution ($s_{3}$) $51$
Number of BinConv blocks ($M$) $3$
Dropout $0.35$
Learning rate $0.001$
Table 11: Number of trainable parameters (in thousands) for each model across datasets.
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
ETTh1 130.368 148.240 209.059 302.071 430.001 637.663
ETTh2 18.624 33.616 414.983 115.815 430.001 637.663
ETTm1 130.368 1347.872 264.263 297.079 430.001 637.663
ETTm2 18.624 1347.872 412.459 70.199 430.001 637.663
Illness 2.664 57.836 588.039 69.623 430.001 103.063
Table 12: Model rankings by NMAE and CRPS (lower is better). Rank 1 = best performance. Averages are computed across all datasets.
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
NMAE Rank
ETTh1 3 2 5 4 6 1
ETTh2 3 1 6 4 5 2
ETTm1 1 2 5 6 3 4
ETTm2 3 1 6 5 2 4
Illness 5 3 1 2 6 4
Average 3.0 1.8 4.8 4.2 4.4 3.0
CRPS Rank
ETTh1 4 3 5 1 6 2
ETTh2 5 2 6 3 4 1
ETTm1 2 3 5 6 1 4
ETTm2 4 3 6 5 1 2
Illness 5 3 2 1 6 4
Average 4.0 2.8 4.8 3.2 3.6 2.6
Appendix B Qualitative Analysis
B.1 The Ability to Capture the Trend
In Section 6.1, we mentioned that replacing the convolutional layer with a fully connected layer as the final layer results in the model’s inability to forecast values outside the training range. This occurs because, unlike convolutional layers, fully connected layers do not share weights. As a result, each component of the output vector is associated with its own set of learnable parameters that influence only that component.
A direct consequence of this is that if the last $K$ components of the labels during training are always zero, the network can trivially learn to predict zeros for those components, thereby failing to extrapolate any upward or downward trends.
To illustrate this, we generated a synthetic dataset with a linear trend:
$s_{t}=(100+1.5\cdot t)(1+\sigma_{t}),\quad t=0,1,\dots,T-1,$
where $\sigma_{t}\sim\mathcal{N}(0,10^{-4})$. We set $T=144$, using the first 120 values for training and the remaining 24 for testing. We then trained both the original BinConv and a variant with a fully connected output layer (BinConv w/ FC), and forecasted the final 48 values. The hyperparameters used for BinConv are provided in Table 13. Context length was set to $72$, and the prediction horizon to $24$. The BinConv w/ FC variant was constructed as described in Section 6.1. For simplicity, during forecasting, we did not sample from the output distribution but instead used the BinConv Argmax procedure from Section 4.2.
Figure 5 shows the forecast for the last 48 values alongside the ground truth. As the test phase begins, BinConv w/ FC outputs a constant value corresponding to the maximum it observed during training, failing to capture the trend. In contrast, BinConv successfully extrapolates the linear trend and achieves accurate forecasts.
Figure 5: Forecast comparison of BinConv and BinConv w/ FC on synthetic linear trend data. The vertical line marks the start of the test data. BinConv w/ FC fails to forecast values beyond those seen during training.
Table 13: Model Parameters for Synthetic Datasets
Parameter Value
Learning Rate (lr) $0.1$
Dropout $0.0$
Number of Bins $500$
Minimum bin value ($b_{0}$) -2
Maximum bin value ($b_{D}$) 2
Number of Blocks ($K$) $1$
Kernel Size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel Size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel Size of final 1D convolution ($s_{3}$) $51$
Number of Filters (2D) $72$
Number of Filters (1D) $72$
B.2 The Ability to Capture the Seasonality
In this section, we illustrate that BinConv is capable of capturing seasonality. To demonstrate this, we generate a synthetic dataset over discrete time steps $t=0,1,\dots,T-1$. The data follows a sinusoidal seasonal pattern with additive Gaussian noise:
$s_{t}=100+A\cdot\sin\left(\frac{2\pi t}{P}\right)+\varepsilon_{t},\quad t=0,1,\dots,T-1,$
where $A=10$, $P=12$, $\varepsilon_{t}\sim\mathcal{N}(0,1)$, and $T=144$. The model hyperparameters are reported in Table 13. The context length is set to 72 and the prediction horizon to 24. The model was trained on the first 120 data values and used to forecast the last 48. The first 24 test values were seen during training, while the final 24 values were not.
Figure 6 presents the forecast results, which demonstrate that BinConv successfully captures the seasonal structure of the data.
Figure 6: Comparison of forecasting methods against observed values. The vertical line indicates the start of the test data.
Appendix C Evaluation Metrics
In our experiments, we employ the normalized mean absolute error (NMAE) for evaluating point forecasts, and the continuous ranked probability score (CRPS) for evaluating probabilistic forecasts. These metrics are widely adopted in time series forecasting research [zhang2024probts, chen2020probabilistic, ansari2024chronos, rasul2021multivariate].
NMAE is defined as:
$\mathrm{NMAE}=\frac{\sum_{k=1}^{K}\sum_{t=1}^{T}\left|x_{t}^{k}-\hat{x}_{t}^{k}\right|}{\sum_{k=1}^{K}\sum_{t=1}^{T}\left|x_{t}^{k}\right|},$
where $K$ denotes the dimensionality of the multivariate time series (and is $1$ in the univariate case), $T$ is the prediction length (i.e., forecasting horizon), $x_{t}^{k}$ is the observed target value, and $\hat{x}_{t}^{k}$ is the predicted value.
CRPS is given by:
[…]
where $\Lambda_{\alpha}(q,z)$ is the quantile loss function at quantile level $\alpha$, defined as:
$\Lambda_{\alpha}(q,z):=\left(\alpha-\mathbb{I}{\{z<q\}}\right)(z-q),$
and $\mathbb{I}{\{z<q\}}$ is the indicator function, which equals $1$ if $z<q$ and 00 otherwise.
To approximate CRPS for discrete data, we use 19 quantile levels ranging from $0.05$ to $0.95$ in steps of $0.05$, leading to:
$\operatorname{CRPS}(F^{-1},x)\approx\sum_{i=1}^{19}2\Lambda_{i\times 0.05}\left(F^{-1}(i\times 0.05),x\right).$
To estimate the empirical quantile function $F^{-1}$, we sample 100 forecasts. For probabilistic forecasting models, the point forecast is computed as the average of these sampled forecasts. In contrast, for point forecasting models, all quantiles are presented by the single predicted value. In this case, it can be shown that the NMAE is equivalent to the CRPS, leading to identical values for both metrics.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly outline the proposed BCE representation and BinConv architecture. These contributions are evaluated empirically and their benefits are substantiated through experiments.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed explicitly in the conclusion section, including discretization issues, bin configuration challenges, and the restriction to univariate time series.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper is empirical in nature and does not present theoretical results requiring formal proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Detailed information is provided regarding datasets, training procedures, evaluation metrics, and model configurations, which allows others to reproduce the experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper uses publicly available datasets (e.g., M4, Tourism) and commits to releasing code and instructions for reproducibility in supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies dataset splits, training epochs, evaluation metrics, hyperparameters, and model architecture clearly in both text and tables.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Standard deviation values are reported for timing and metrics tables. Table 3 shows ranges across runs, demonstrating robustness.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: GPU type (NVIDIA A10), memory, and timing per sample are reported in Table 6, allowing estimation of computational cost.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work uses publicly available datasets and does not involve sensitive or personal data. Ethical considerations have been followed.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper focuses on a technical contribution to time series forecasting using novel representation and modeling techniques. The methods are domain-agnostic and do not directly involve applications with immediate societal implications. As such, the work does not raise specific concerns related to societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release high-risk models or data requiring safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets used (e.g., M4, Tourism) are publicly available and properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The BinConv model and BCE encoding are novel contributions of this work. An anonymized zip file containing the implementation is included, and the code will be publicly released to ensure full reproducibility.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The research does not involve human subjects or crowdsourcing.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects are involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: While LLMs were used for minor editing and formatting suggestions, they did not influence core methodology or scientific contributions.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.